### v1

# Attention-Based Reading, Highlighting, and Forecasting of the Limit Order Book

Abstract
Managing high-frequency data in a limit order book (LOB) is a complex task that often exceeds the capabilities of conventional time-series forecasting models. Accurately predicting the entire multi-level LOB, beyond just the mid-price, is essential for understanding high-frequency market dynamics. However, this task is challenging due to the complex interdependencies among compound attributes within each dimension, such as order types, features, and levels. In this study, we explore advanced multidimensional sequence-to-sequence models to forecast the entire multi-level LOB, including order prices and volumes. Our main contribution is the development of a compound multivariate embedding method designed to capture the complex relationships between spatiotemporal features. Empirical results show that our method outperforms other multivariate forecasting methods, achieving the lowest forecasting error while preserving the ordinal structure of the LOB.
keywords:
limit order book , multivariate time-series forecasting , attention mechanism , spatiotemporal embedding
[…]
[inst1]organization=Department of Statistics, Purdue University,city=West Lafayette, postcode=47907, state=IN, country=USA
1 Introduction
A limit order book (LOB) is a digital record of all outstanding buy and sell orders for a particular financial instrument, such as a stock or currency. It contains information about the price and quantity of each order, as well as the time at which the order was placed. The LOB continuously updates as new orders are submitted, executed, or canceled, providing an up-to-the-moment snapshot of market supply and demand. This evolution of orders highlights the distribution of orders across various price levels, providing information on market depth and potential price movements.
However, predictive modeling of LOBs presents considerable challenges due to the high frequency and complexity of the data. The dynamic nature of the LOB, with orders being added, modified, or removed at millisecond intervals, requires rapid processing and analysis of vast amounts of information. Additionally, LOB data has a complex multivariate time-series structure with compound attributes, where levels, types, and features are interrelated and simultaneously associated with each value. For instance, the ordinal structure in multi-level prices highlights the interconnectedness of these attributes. Therefore, accurately capturing the influence of each attribute is essential for effective predictive modeling of the LOB.
Given the complexities involved in predicting the full dynamics of a limit order book (LOB), many researchers have narrowed their focus to forecasting the mid-price, which is the average of the best bid and ask prices. Various studies have utilized different models for this task, including regression-based approaches (Ntakaris et al., 2018), support vector machines (Kercheval and Zhang, 2015), and deep neural networks (Sirignano, 2019; Zhang et al., 2019) to predict future mid-price movements. While mid-price serves as a useful indicator of overall market trends, Figure 1 highlights that the LOB can exhibit different features such as order depths at different price levels and bid-ask spreads (BAS), even when the mid-price remains the same. Therefore, relying solely on mid-price analysis may be insufficient when more detailed depth information is needed, such as in optimal execution problems, where a comprehensive understanding of the full LOB structure is essential.
Figure 1: Three different LOB snapshots with the same mid-price: The three limit order book snapshots share the same mid-price, with the first and second snapshots showing varying levels of depth available, while the third snapshot features a larger bid-ask spread.
Attention-based models like Transformer (Vaswani et al., 2017) and BERT (Devlin et al., 2019) excel at sequence-to-sequence prediction by assigning varying weights to different parts of the data, enabling them to focus on the most relevant information for more accurate predictions. These models have proven highly effective in tasks such as natural language processing (Yang et al., 2021; Galassi et al., 2020; Cui et al., 2019), speech recognition (Zeyer et al., 2018; Chorowski et al., 2015; Bahdanau et al., 2016), and time series forecasting (Shih et al., 2019; Fan et al., 2019; Du et al., 2020). However, their application to limit order book forecasting remains limited, with notable examples including Arroyo et al. (2024) in survival analysis, Liu et al. (2024) using vision transformers, and Wu et al. (2024) employing wavelet decomposition.
In this study, we propose employing attention-based models, particularly those introduced in Spacetimeformer (Grigsby et al., 2022), to address the unique challenges posed by the spatiotemporal forecasting of limit order book. Our approach builds upon the ideas presented in this work and applies them specifically to the task of predicting limit order book movements. Our embedding approach integrates spatiotemporal features while preserving the structure of the LOB, leading to superior predictive accuracy compared to traditional forecasting methods. These advancements demonstrate the potential of attention-based models to enhance predictive performance in high-frequency trading environments.
2 Data and method
2.1 Data structure and preparation
We utilize the LOBSTER dataset (Huang and Polak, 2011), which offers real-world LOB data, including both message and order book information for selected stocks. We specifically focus on five major technology companies: AAPL, GOOG, INTC, MSFT, and AMZN. The sample covers the trading day of June 21, 2012, from the market opening at 9:30 am to the closing at 4:00 pm, capturing between 300,000 to 600,000 snapshots per stock, all recorded with millisecond time resolution at Level-5 of the order book. The order book file tracks the evolution of the LOB, while the message file logs the events that trigger updates, with timestamps ranging from milliseconds to nanoseconds.
The order book consists of two types of orders: bids (to buy) and asks (to sell). An LOB is continuously updated with every change in these orders, including new submissions, cancellations, deletions, or executions. Each snapshot of the LOB data is represented as $\vec{X_{i}}$ for $i\in\{1,\dots,I\}$. Both bid and ask orders are captured as vectors containing the timestamp and price-volume pairs for each level $k\in{1,\dots,K}$. Specifically, for the $i^{\text{th}}$ snapshot, $\vec{X_{i}}=\left[t_{i},(p_{1i}^{b},v_{1i}^{b}),\dots,(p_{Ki}^{b},v_{Ki}^{b}),(p_{1i}^{a},v_{1i}^{a}),\dots,(p_{Ki}^{a},v_{Ki}^{a})\right]$, where $t_{i}$ is the timestamp, $p_{ki}^{b/a}$ represents the level-$k$ bid/ask price, and $v_{ki}^{b/a}$ represents the level-$k$ bid/ask volume. To maintain consistency, we fix the number of levels, $K$, at 5 for both bid and ask orders in our experiments.
To ensure consistency and facilitate analysis, we standardized the timestamps to equal intervals of 5 seconds during data cleansing. For example, this results in approximately 4,680 snapshots per trading day. We also concatenated the data from all five stocks, creating a dataset with 100 dimensions, including five levels of bid and ask prices and volumes for each stock. The dataset was divided into training, validation, and testing sequences in a 6:2:2 ratio. For prediction, a context sequence of 120 time steps (equivalent to 10 minutes) was used to output a target sequence of 24 time steps (equivalent to 2 minutes).
2.2 Stationary transformation and scaling
Figure 2: Input encoding pipeline for limit order book forecasting. (1) Level-5 LOB sequences are split into 10-minute context points and 2-minute target points. (2) Time, variable, and context embeddings are created to structure the input sequence. (3) The final LOB input sequence is generated by integrating these embedding layers.
The input encoding pipeline for LOB prediction involves pre-processing the LOB data to be fed into a forecasting model (Figure 2). This step involves extracting relevant features from the LOB data, such as the price, volume, and time of each order. However, non-stationary dimensions in the data can hinder learning by causing distributional shifts between the training and testing periods, a common challenge when dealing with long sequences of time series data. For example, multi-level prices often exhibit non-stationarity even within a single trading day.
In our experiment, we employ percent-change transformation of prices to facilitate training. The percent change transformation is applied to multivariate price sequences at all levels using the following equation:
${p}_{k,i}^{\text{perc}}=\frac{{p}_{k,i}-{p}_{k,i-1}}{{p}_{k,i-1}}$ (1)
Figure 3 illustrates that the transformed series exhibits greater stationarity in price movements compared to the raw series. During training, we also apply min-max scaling to price and volume on a variable-wise basis to ensure that variables with large fluctuations do not disproportionately influence the model. The min-max scaling transformation for price and volume is defined as:
${p}_{k,i}^{\text{scaled}}=\frac{{p}_{k,i}^{\text{perc}}-\min_{i}({p}_{k,i}^{\text{perc}})}{\max_{i}({p}_{k,i}^{\text{perc}})-\min_{i}({p}_{k,i}^{\text{perc}})},\quad{v}_{k,i}^{\text{scaled}}=\frac{{v}_{k,i}-\min_{i}({v}_{k,i})}{\max_{i}({v}_{k,i})-\min_{i}({v}_{k,i})}$ (2)
Figure 3: Price movement of a limit order book: raw time series (left) and percent-change transformed time series (right).
2.3 Compound multivariate embedding
The embedding process encodes time, values, variables, and context-target information. We employ the Time2Vec layer (Kazemi et al., 2019) to transform timestamps into frequency embeddings, which are then passed through a feed-forward layer to capture temporal patterns and data sequence. Additionally, a binary context-target embedding identifies whether a value is from the context or is to be forecasted, enabling the model to distinguish between observed and future data points.
Our data structure incorporates compound attributes such as multiple levels (level-$k$), types (bid or ask), features (price or volume), and stocks (tickers). To effectively capture dependencies among these attributes, we developed a compound multivariate embedding method inspired by the spacetimeformer approach (Grigsby et al., 2022). Unlike the original spacetimeformer, which assigns embedding integers to each variable separately, our method assigns embeddings to each attribute. Then, we combine and scale the multi-level embeddings. As illustrated in Figure 4, each attribute is represented by its own embedding layer. For instance, one layer might encode whether the data is a bid or ask, another might encode the order level (e.g., level-1, level-2), and another might encode the feature type (price or volume).
By integrating and scaling these multiple embedding layers, our approach captures not only temporal and spatial patterns but also the complex relationships between different attributes within the LOB data. This method also reduces the number of parameters in the embedding layer while effectively incorporating interdependencies across attributes. The embedded sequence is processed using an attention-based encoder-decoder architecture (Figure 5). The encoder employs self-attention modules to capture both local and global contexts, while the decoder uses masked attention to prevent access to future information during prediction. The final feed-forward layer transforms the output back into the original LOB sequence format.
The attention mechanism (Vaswani et al., 2017) updates token representations and re-normalizes the resulting matrix:
$\text{Attention}(Q,K,V)\in\mathbbm{R}^{L_{x}\times d}=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)V,$ (3)
where query vectors $Q=W^{Q}X$, key vectors $K=W^{K}Z$, and value vectors $V=W^{V}Z$ are derived from learned parameters. For a sequence of length $L$ and dimension $d$, the self-attention matrix outputs an $(L\times d)^{2}$ matrix, enabling assessment of both temporal and inter-variable dependencies. In our experiment, we use the Performer method (Choromanski et al., 2020) to enhance computational efficiency, which approximates attention with linear space and time complexity.
Figure 4: Input encoding pipeline for limit order books: spatiotemporal (A) and compound multivariate (B) embedding.
Figure 5: The Spatiotemporal attention architecture applied to the sequence of limit order books. The darker shades indicate higher attention scores. This approach is based on the work of Grigsby et al. (2022).
2.4 Training
Forecasting loss
For our evaluation, we measure the difference between predicted and actual limit orders using mean squared error (MSE) and mean absolute error (MAE) as our primary metrics. The forecasting loss specifically focuses on minimizing the MSE of level-5 price and volume predictions within our sample dataset:
Forecasting $\text{Forecasting Loss}=\sum_{i,j,k}({p}_{k,i}^{j}-\hat{p}_{k,i}^{j})^{2}+\sum_{i,j,k}({v}_{k,i}^{j}-\hat{v}_{k,i}^{j})^{2}.$ (4)
MSE is chosen because of its effectiveness in addressing multivariate forecasting tasks, particularly in transformer-based models, and its ability to reduce significant discrepancies in predictions. Although metrics like MAE are recorded for reference, they are not used as loss functions.
Structural regularizer
To ensure that our model incorporates the multi-level limit order book (LOB) structures during the training process, we introduce a structural regularizer. For instance, the level-1 ask (bid) price is defined as the lowest (highest) offer price available from sellers (buyers) quoting a security, while level-2 prices represent the second-best offers, and so forth. Let $p_{ki}^{a}$ ($p_{ki}^{b}$) represent the level-$k$ ask (bid) price for the $i$-th snapshot, where $k$ denotes the order level. For any fixed $i$-th LOB snapshot, the multi-level price structure is given by:
$p_{k_{1}i}^{a}<p_{k_{2}i}^{a}$ for all $\quad\text{for all }k_{1}<k_{2},$ (5)
$p_{k_{1}i}^{b}<p_{k_{2}i}^{a}$ for $\quad\text{for }k_{1}=k_{2}=1,$ (6)
$p_{k_{1}i}^{b}>p_{k_{2}i}^{b}$ for all $\quad\text{for all }k_{1}<k_{2}$ (7)
To nudge our network outputs adhere to these ordinal price structures, we add a regularizing term to capture this restriction at each snapshot-$i$:
$\text{Structure loss}_{i}=\sum_{k=1}^{K-1}\left(\text{ReLU}(\hat{p}_{k,i}^{a}-\hat{p}_{k+1,i}^{a})+\text{ReLU}(\hat{p}_{k+1,i}^{b}-\hat{p}_{k,i}^{b})\right)+\text{ReLU}(\hat{p}_{1,i}^{b}-\hat{p}_{1,i}^{a}),$ (8)
where $p_{k,i}^{j}$ refers to the actual price, $\hat{p}_{k,i}^{j}$ refers to the predicted price, and the ReLU (Rectified Linear Unit) function is defined as $\text{ReLU}(x)=\max(0,x)$. This function ensures that only positive differences contribute to the penalty, thus enforcing the ordinal structure by penalizing any violation where the predicted lower-level ask price is not less than the higher-level ask price, or where the predicted lower-level bid price is not greater than the higher-level bid price.
The total loss function is defined with a regularization weight $w_{o}\geq 0$, which determines the strength of the regularization:
Loss=Forecasting $\text{Loss}=\text{Forecasting loss}+w_{o}\cdot\sum_{i=1}^{I}\text{Structure loss}_{i},$ (9)
This function penalizes predictions that break the ordinal structure of price levels while also minimizing the forecasting error, thereby ensuring that the correct multi-level sequences are maintained during training. We fixed $w_{o}$ at 0.01 in our experiment for balancing the scale of forecasting and structure losses.
For training, the learning rate decay is managed with a decay factor of 0.8 and 1000 warmup steps. Additionally, the model employed seasonal decomposition and reversible normalization (Kim et al., 2021) techniques to enhance performance. We use three-head attention modules for training the attention-based models. We monitored the validation loss across epochs and halted the process if no improvement was observed after 10 consecutive epochs. For testing, we selected the trained model weights associated with the lowest validation total loss. The training was conducted on Google Colab using a single A100 GPU.
3 Results
Table 1: Prediction results for sample test data: Mean Squared Error (MSE) and Mean Absolute Error (MAE) are computed for each method across selected dimensions. The mid-price is determined by averaging the best bid and ask prices. Loss metrics are provided for price, volume, and overall performance on level-5 limit order book (LOB) data. Structure loss quantifies the violations of the ordinal structure, while total loss represents the weighted sum of the forecasting loss. All values are calculated after scaling.
Time series models Attention-based models
Linear LSTM Temporal Spacetime Compound (ours)
[…]
    Price metrics refer to the aggregated loss for bid and ask prices across all five levels.
    Volume metrics refer to the aggregated loss for bid and ask volumes across all five levels.
    Entire refers to the sum of losses from all five levels of prices and volumes.
We present the prediction results for sample test data, with a focus on various performance metrics discussed in Section 2.4. Our experiments compare various sequence-to-sequence prediction models of prior work, including a linear autoregressive model (Linear), a standard LSTM network (Schmidhuber et al., 1997) (LSTM), a transformer with temporal attention (Temporal) as introduced in Informer (Zhou et al., 2021), and a spacetimeformer model (Spacetime) that incorporates both spatial and temporal attention (Grigsby et al., 2022). We also evaluate the performance of our attention-based model with compound multivariate embedding (Compound) in comparison to the other techniques.
As shown in Table 1, our compound multivariate embedding method consistently ranked first or second among both time series and attention-based models, achieving the lowest forecasting error across the entire dataset. Notably, the compound multivariate embedding model significantly outperformed all others in structure loss—by a ratio of 0.2 to 0.6 times smaller—indicating far fewer violations of the ordinal structure in its predictions. This highlights the model’s superior ability to capture structural dependencies through the compound multivariate embedding process.
Figure 6 and Figure 7 provide visual comparisons of the actual versus predicted price and volume movements for the level-5 LOB data. We selected the first window of the test period for comparison. In Figure 6, the price forecasting results demonstrate that the compound multivariate embedding method (B) closely tracks the downward trend of actual price movements (A), offering more accurate predictions than the temporal (C) and spatiotemporal (D) embedding methods. While both our approach and the spatiotemporal embedding effectively capture the overall ordinal structure of multi-level prices, the temporal embedding fails to preserve the ordinal relationships between dimensions. This aligns with the concepts of the embedding methods, as the temporal embedding processes multiple time series without incorporating variable embeddings, resulting in structural violations. Similarly, in Figure 7, the volume forecasting graphs show that the compound multivariate embedding method (B) yields overall volume distributions that more closely match the actual order sizes (A), outperforming the other attention-based models (C and D).
These empirical results underscore the effectiveness of the compound multivariate embedding method in capturing the intricate dependencies within the LOB data, leading to more accurate and reliable forecasts for both price and volume.
Figure 6: Price forecasting of a limit order book for AMZN stock: The graphs predict level-5 prices 2 minutes ahead, based on the preceding 10 minutes of context, with the prediction period starting on the right side of the dotted line. Panel A (top left) shows the actual price movements, while panels B, C, and D display the predicted price movements using three attention-based methods: compound multivariate embedding (B), temporal embedding (C), and spatiotemporal embedding (D). Multi-level prices, excluding the best prices, are represented in sequential grey lines.
Figure 7: Volume forecasting of a limit order book snapshot for AMZN stock: The graphs forecast level-5 volumes for the first snapshot of the testing period based on the preceding 10 minutes of context. Panel A shows the actual order sizes, while panels B, C, and D display the predicted volume snapshots using three attention-based methods: compound multivariate embedding (B), temporal embedding (C), and spatiotemporal embedding (D).
4 Conclusions
In this study, we addressed the challenge of forecasting multi-level limit order book (LOB) data, a task that often exceeds the capabilities of conventional time-series forecasting models. Our primary contribution is the development and implementation of a compound multivariate embedding method within advanced sequence-to-sequence models. This method effectively captures the complex interdependencies among various LOB attributes, including order types, features, and levels. The empirical results demonstrate that our approach consistently outperforms existing multivariate forecasting methods, achieving the lowest forecasting errors while maintaining the ordinal structure of the LOB. These findings highlight the effectiveness of our method for financial time-series forecasting, particularly in the context of high-frequency trading. As a direction for future work, our method could be extended to address optimal execution problems, with further optimization for real-time forecasting scenarios to enhance its practical applicability in dynamic trading environments.
Conflict of interest
The authors have no conflict of interest to disclose.
Data availability
The authors do not have permission to share data.